Please help me step by step and as clearly as possible PLEASE, Please treat me like an absolute idiot and list out every step by step please. I got this home work that says the following:

Find all Jordan Canonical forms of a 3x3 matrix A with. ( what does it mean when saying I can assume that the) diagonal entries are in increasing order.

det(A – xIq) = (2 – x)(1 – x)2.
Answer #1

Jordan Canonical form & If de matrix Anonhas characteristic equation, polynomial CA (A) = (x-A) (andjo then, Jordan will beIn our question : det (A-x Ia) = (2-x3 (1-xp. 5) Characteristic, equation (co () = (dame)? (2x-1) ? polynomial since minimalEsplanation about [ma (dl), (mA, - (2-2)! (2-1) by first sule (sule Opohrich be we mentioned earty), jordan forms will contai